Suppose that the demand for crude oil is given by Qd =85- 0.4P, where Qd is the quantity demanded in millions of barrels per day and P is the price per barrel in dollars. Suppose also that the supply of crude oil is given by Qs= 55 + 0.6P. a) Calculate equilibrium price and quantity in this Market b) Graph the demand curve and the supply curve c) Calculate the demand for crude oil and the supply for crude oil if the market price is $ 15.00 per barrel. Explain your answer. d) Calculate the demand for crude oil and the supply for crude oil if the market price is $50.00 per barrel. Explain your answer. 2. What type of relationship exists between the firm’s average physical product curve and the average cost curve in the short- run


1
Expert's answer
2021-10-03T14:14:48-0400

Given

"Qd=85-0.4p\\Qs=55+0.6p"

.a)

"Qd=Qs\\"

"85-0.4p=55+0.6p\\85-55=0.4p+0.6p\\p=30\\Q=55+06(30)\\Q=73"

b)


c)

"Qd=85-0.4p\\=85-0.4(15)\\=79\\Qs=55+0.6p\\=55+0.6(15)\\=64"

d)

"Qd=85-0.4p\\=85-0.4(50)\\=65\\Qs=55+0.6p\\=55+0.6(50)\\=85"

2. Firm’s average physical product curve is horizontal while the average cost curve is U-shaped in the short- run